OpenStudy (anonymous):

Need help please! The endpoint of the diameter of circle have coordinates at (x+2 , y+3) and (3x-6) , (5y-7).Find the coordinates of the center of the circle.

OpenStudy (anonymous):

use the midpoint formula

OpenStudy (anonymous):

[(x+2)+(3x-6)]/2=2x-2

OpenStudy (anonymous):

[(x+2)+(3x-6)]/2,[(y+3)+(5y-7)]/2

OpenStudy (anonymous):

y=[(y+3)+(5y-7)]/2=2/3?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

y=2/3,x=1?

OpenStudy (anonymous):

the coordinates of the center are (2x-2),(3y-2)

OpenStudy (anonymous):

is it given that x = 1, y = 2/3?

OpenStudy (anonymous):

[(x+2)+(3x-6)]/2 =(4x-4)/2 =2x-2 [(y+3)+(5y-7)]/2 =(6y-4)/2 = 3y-2 stop from here , on more father?

OpenStudy (anonymous):

I think you can use the distance formula. the distance from the center to the two end points are the same. so equate both the distances using the distance formula, but that wont solve it. you need two equations to solve it.

OpenStudy (anonymous):

I know, if no given x, y than I don't have to solve ,Just write two equations only right ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you are welcome